OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by the curves y=4+sec(3x) ,between -pi/9 and pi/9 , y=6 about the line y-4

OpenStudy (anonymous):

\[\pi\int_{-\frac{\pi}{9}}^{\frac{\pi}{9}}(4+sec(3x))dx\]

OpenStudy (anonymous):

is that your final answer?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

about the line y = 4?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\pi\int_{-\frac{\pi}{9}}^{\frac{\pi}{9}}((4+sec(3x))^2-(4^2))dx\]

OpenStudy (anonymous):

i got all those steps i need to see what your final answer is cos i keep getting my final answer wrong